A car that weighs 7,000 N is accelerated from rest to a velocity of 25 m/s eastward by a force of 1,000 N. (a) What was the change in momentum of the car? (b) How long did the torce a to change the momentum of the car?



1
Expert's answer
2022-06-13T08:23:02-0400

Given:

"mg=7000\:\rm N"

"F=1000\:\rm N"

"\Delta v=25\:\rm m/s"


(a)

"\Delta p=m\Delta v\\
=7000/9.8*25=17500\:\rm kg\cdot m/s"

(b)

"F=\frac{\Delta v}{\Delta t}"

"\Delta t=\frac{\Delta v}{F}=\frac{17500}{1000}=17.5\:\rm s"